OpenStudy (anonymous):

Find (fraction): 5/9 - 3/4 =

OpenStudy (anonymous):

7/36 ?

OpenStudy (anonymous):

-7/36

OpenStudy (anonymous):

20 - 27 -7 5/9 - 3/4 = ----------- = ------ 36 36

OpenStudy (anonymous):

Many Thanks is 2/9 - 1/6 + 1/2 = 5/9 having it simplified?. Want to check my answer :)

OpenStudy (anonymous):

ya. simplified is lowest term. and correct

OpenStudy (anonymous):

12 - 9 + 27 30 5 2/9 - 1/6 + 1/2 = ------------ = ----- = ------ 54 54 9 so yr ans is correct....☺

OpenStudy (anonymous):

Thank You :)!